OpenStudy (anonymous):

Find cos Θ/2 when 270º<Θ<360º and cosΘ is defined as indicated. cos Θ =√7/4

OpenStudy (misty1212):

\[\cos(\frac{\theta}{2})=\pm\sqrt{\frac{1+\cos(\theta)}{2}}\]

OpenStudy (misty1212):

\[\cos(\frac{\theta}{2})=\pm\sqrt{\frac{1+\frac{\sqrt7}{4}}{2}}\]

OpenStudy (misty1212):

use the \(-\) not the \(+\) because half an angle in quadrant 4 is in quadrant 2, and in quadrant 2 cosine is negative

OpenStudy (anonymous):

is it \[-\sqrt{4-\sqrt{7}}/4+\sqrt{7}\]

OpenStudy (anonymous):

@misty1212